# An $H^{1}$ multiplier theorem on anisotropic Hardy sapces

### v1

@@ 4. Further results
Using the above method, few more results can be obtained. Under the isotropic settings, the following theorem was announced in [8].
@@
Using Proposition 3.2 and Hölder’s inequality, we know that for $k\leq-(l+N)$,
@@
Now we apply $f=g\cdot\widehat{\chi_{k,\lambda}}$ to (4.1), it follows that $\left(\int|\widehat{g}\ast\chi_{k,\lambda}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim b^{\frac{k}{r}}\|g\|_{r}$. By duality, we know that for all $h\in L^{p^{\prime}}(\mathrm{d}\mu)$,
@@
By Fubini’s theorem, it follows that
@@
We begin with a square function characterization of $H^{1}$. Let $\eta$ be a Schwartz function on $\mathbb{R}^{d}$ whose Fourier transform is nonnegative, supported in $(A^{\ast})^{2}(\Delta^{\ast})\setminus(A^{\ast})^{-1}(\Delta^{\ast})$, equal to $1$ on $(A^{\ast})(\Delta^{\ast})\setminus\Delta^{\ast}$. We define $\Psi$ via the Fourier transform:
@@
Therefore,
$\left|\widehat{f}(\xi)\right|\leq\sum_{l-2\leq j\leq l+2}\left\|\widehat{\Psi_{j}\ast f}\right\|_{\infty}\leq\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}.$
@@
$\lesssim\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}^{p},$
where we used the hypothesis (4.3) and the trivial estimate $\left(\sum_{i=1}^{5}|a_{i}|\right)^{p}\lesssim\sum_{i=1}^{5}|a_{i}|^{p}$. Summing over $l\in\mathbb{Z}$ implies that
$\left(\int|\widehat{f}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{p}\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}},$
where we used the condition $p\geq 2$ in the last inequality. Finally, the integral form of Minkowski’s inequality yields
@@
Therefore, the $H^{1}$ norm of $\Psi_{-l}$ is independent of $l$. However, from the construction of $\Psi$, we know that

### v3

@@ 5. Further results
Under the isotropic settings, the following theorem was announced in [8].
@@
Using Proposition 4.2 and Hölder’s inequality, for $k\leq-(l+N)$,
@@
We apply $f=g\cdot\widehat{\chi_{k,\lambda}}$ to (5.1), it follows that $\left(\int|\widehat{g}\ast\chi_{k,\lambda}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim b^{\frac{k}{r}}\|g\|_{r}$. By duality, for all $h\in L^{p^{\prime}}(\mathrm{d}\mu)$,
@@
By Fubini’s theorem,
@@
We begin with a square function characterization of $H^{1}$. Let $\eta$ be a Schwartz function on $\mathbb{R}^{d}$ whose Fourier transform is nonnegative, supported in $(A^{\ast})^{2}(\Delta^{\ast})\setminus(A^{\ast})^{-1}(\Delta^{\ast})$, equal to $1$ on $(A^{\ast})(\Delta^{\ast})\setminus\Delta^{\ast}$. We define $\Psi$ via the Fourier transform,
@@
Taking supremum over $\xi$,
$\sup_{\xi\in(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})}\left|\widehat{f}(\xi)\right|\leq\sum_{l-2\leq j\leq l+2}\left\|\widehat{\Psi_{j}\ast f}\right\|_{\infty}\leq\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}.$
@@
$\lesssim\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}^{p}.$
where we used the hypothesis (5.3) and the trivial estimate $\left(\sum_{i=1}^{5}|a_{i}|\right)^{p}\lesssim\sum_{i=1}^{5}|a_{i}|^{p}$. Summing over $l\in\mathbb{Z}$, and use the condition $p\geq 2$,
$\left(\int|\widehat{f}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{p}\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}}.$
Finally, the integral form of Minkowski’s inequality yields
@@
Therefore, the $H^{1}$ norm of $\Psi_{-l}$ is independent of $l$. From the construction of $\Psi$,